(Log75 +Log9+Log5) ÷ (Log5 + Log45)


1
Expert's answer
2021-04-27T12:40:14-0400

"(Log75 +Log9+Log5) ÷ (Log5 + Log45)={Log75*9*5 \over Log5*45}={Log25*3*9*5 \over Log5*9*5}={Log5^3*3^3 \over Log5^2*3^2}={Log15^3 \over Log15^2}={3 \over 2} =1.5"